1.what will be the diagonal of a square whose perimeter is 60m?
2.The sides of a rectangular field of area 726 sq meters are in the ratio 3:2 . find the length of its diagonal correct up to one decimal place ?

Answer by jim_thompson5910(35256)

I'll do the first one to get you started

# 1 what will be the diagonal of a square whose perimeter is 60m?

Side Length = 60/4 = 15

a^2 + b^2 = c^2

15^2 + 15^2 = c^2

225 + 225 = c^2

c^2 = 2*225

c = sqrt(2*225)

c = sqrt(2)*sqrt(225)

c = sqrt(2)*15

c = 15*sqrt(2) <---- Exact length

c = 21.2132 <---- Approximate length